# EDVD-LLaMA: Explainable Deepfake Video Detection via Multimodal Large Language Model Reasoning

### v1

@@ Abstract
The rapid development of deepfake video technology has not only facilitated artistic creation but also made it easier to spread misinformation, which is increasingly difficult to identify. Traditional deepfake video detection (DVD) methods face issues such as a lack of transparency in their principles and insufficient generalization capabilities to cope with evolving forgery techniques. This highlights an urgent need for detectors that can identify forged content and provide verifiable reasoning explanations. This paper proposes the explainable deepfake video detection (EDVD) task and designs the EDVD-LLaMA multimodal, a large language model (MLLM) reasoning framework, which provides traceable reasoning processes alongside accurate detection results and trustworthy explanations. Our approach first incorporates a Spatio-Temporal Subtle Information Tokenization (ST-SIT) to extract and fuse global and local cross-frame deepfake features, providing rich spatio-temporal semantic information input for MLLM reasoning. Second, we construct a Fine-grained Multimodal Chain-of-Thought (Fg-MCoT) mechanism, which introduces facial feature data as hard constraints during the reasoning process to achieve pixel-level spatio-temporal video localization, suppress hallucinated outputs, and enhance the reliability of the chain of thought. In addition, we build an Explainable Reasoning FF++ benchmark dataset (ER-FF++set), leveraging structured data to annotate videos and ensure quality control, thereby supporting dual supervision for reasoning and detection. Extensive experiments demonstrate that EDVD-LLaMA achieves outstanding performance and robustness in terms of detection accuracy, explainability, and its ability to handle cross-forgery methods and cross-dataset scenarios. Compared to previous DVD methods, it provides a more explainable and superior solution. The source code and dataset will be publicly available.
@@ I Introduction
With the increasing popularity of Artificial Intelligence Generated Content (AIGC), deepfake media are easily created using publicly available AIGC video editing tools [1, 2, 3]. These techniques can replace faces in videos or alter facial expressions and movements, making such manipulations difficult to detect without a specialized tool [4]. While deepfake videos can be used for artistic creation and entertainment, they have serious incidents such as financial fraud, identity theft, and the fabrication of evidence to manipulate public opinion, posing significant information security risks [5, 6, 7]. Therefore, developing advanced methods for Deepfake Video Detection (DVD) has become critically important. In this context, the task of DVD aims to determine whether a video is a deepfake. DVD has broad applications in daily work and life, such as verifying the authenticity of individuals in video calls, filtering out fake videos on social media, and preventing the spread of malicious video content.
@@
Recent DVD methods have achieved impressive performance by leveraging carefully designed network architectures, sophisticated constraints, and domain-biased pretraining strategies [8, 9]. However, existing DVD approaches face two key challenges that limit their generalization and practicality [10] (see Fig. 2(a)). First, most DVD methods are black-box models that only output a binary classification result and a confidence score for a video, without revealing the basis of their decisions to users. Because these methods lack interpretable decision rationales, manual re-examination is still required. Second, in the real world, video deepfake techniques are diverse and continuously evolving; existing DVD methods lack comprehensive generalization ability and cannot ensure satisfactory accuracy when confronted with newly emerging deepfake techniques, thereby significantly undermining their practical value.
@@
Given that multimodal large language models (MLLMs) are pretrained on large-scale and diverse corpora of world knowledge, recent studies on explainable image forgery detection [11, 12, 13, 14, 15] have introduced MLLM architectures, which show great potential for enhancing the explainability of detection results and have achieved strong performance. However, these methods share two inherent limitations: first, they operate at the image level and are difficult to apply directly to video, resulting in the loss of temporal features and hindering the recognition of holistic video-level deepfake behaviors. Furthermore, the lack of cross-frame consistency modeling prevents key video deepfake cues, such as boundary artifacts and abnormal physiological rhythms in micro-expressions, from being captured. Second, MLLMs often generate descriptions directly from an image, which frequently produces severe hallucinations that list forged artifacts absent from the sample and undermine the credibility of the generated explanations. In addition, although HEIE [15] constructs a Chain-of-Thought (CoT) for static images and achieves promising results, its CoT lacks temporal information. It has no structured data as a hard constraint for structured reasoning. This causes its explanations to rely mainly on free-form text, invites semantic hallucination, and limits generalization to complex and temporally sensitive deepfake video recognition.
@@
Motivated by the abovementioned analysis, we propose an Explainable DVD (EDVD) task and a reasoning framework named EDVD-LLaMA, designed to reason and detect deepfake videos while addressing the two significant limitations. As illustrated in Fig. 2 (b), the EDVD task requires models to extract spatio-temporal features from a given video, integrating pixel-level deepfake clues such as edge artifacts and global semantic anomalies like incorrect perspective, to provide clear reasoning principles and ensure reliable accuracy. In the EDVD-LLaMA framework, we first develop a Spatio-Temporal Subtle Information Tokenization (ST-SIT), which comprises the Deepfake Sniffing Encoder (DSEncoder) for extracting local deepfake features. Meanwhile, SigLiP is utilized to extract global video features, followed by the Compact Visual Connector (CVC) for spatio-temporal compression and fusion. Second, we construct a Fine-grained Multimodal Chain-of-Thought (Fg-MCoT) framework based on LLM [16], which incorporates facial refinement metrics (e.g., landmarks variations) as structured constraints to reduce hallucinations[17] and improve reliability. Additionally, we inject structured facial data into the existing deepfake video datasets [18] and leverage pretrained models [19] to construct question-video-rationale-answer reasoning quadruples, thereby creating an Explainable Reasoning FF++ Dataset (ER-FF++set). Using this dataset, we further fine-tune our EDVD-LLaMA to equip it with comprehensive analytical capabilities, fine-grained reasoning abilities, and accurate deepfake detection capabilities. Our main contributions are summarized as follows:
@@ II Related Work
Video forgery detection. Yan et al. [20] propose video-level blending data (VB) and spatiotemporal adapters (StA) to improve the generalization of deepfake video detection, and for the first time simulate the ”Facial Feature Drift” forgery artifact. Their method balances spatial and temporal information, but still has room for improvement in handling multimodal forgeries and complex real-world scenarios. The PSO-EfficientNet-GRU method [21] leverages particle swarm optimization in conjunction with deep learning models to significantly enhance the accuracy of video forgery detection. UFCC [22] employs content consistency analysis, offering a unified forensic framework that facilitates more stable identification of tampered regions in images and videos. MSIDSnet[23] employs multi-scale fusion and a visual Transformer model to achieve robust detection across videos of varying quality. Overall, these methods enhance the performance of video forgery detection from perspectives such as spatiotemporal features, content consistency, multimodal fusion, and model optimization. However, they still face challenges in generalization ability and adaptation to complex real-world environments.
@@
Explainable image forgery detection. Traditional interpretable deepfake detection methods primarily construct explainable networks through model design [24] or generation operations [25]. For example, DFGNN [26] applies interpretable GNNs to the deepfake detection task. Although these methods enhance the interpretability of the models, they are still unable to describe the reasoning in natural language. In contrast, recent researchers have utilized MLLMs to construct question-answering tasks [11, 13, 12], providing human-understandable explanations for traditional binary classifiers. For example, CSRDD [11] constructs a transferable model based on BLIP, enhancing the generalization ability and interpretability of CNN detectors through multimodal feature fusion. FFAA [13] redefines deepfake image detection as an open-world visual question answering (VQA) task. They propose a Multi-Answer Intelligent Decision System (MIDS), construct an “image description–forensics–decision” triplet dataset, and utilize LLMs combined with multi-hypothesis prompts to generate reasoning explanations for decisions, demonstrating robustness to unknown forgery techniques. FakeShield [12] performs image forgery detection and localization (IFDL) guided by domain labels. A LoRA-tuned MLLM combines tampering classification labels, textual localization rationales, and a segmentation model to achieve synchronous outputs of detection, localization, and supporting evidence, providing interpretable semantic or physical clues and thus improving cross-domain generalization. Although these pioneering works achieve interpretable forgery analysis for single images by integrating visual and language reasoning, they are still limited to static images or facial regions.
@@
Multimodal large language model. With the rapid advancement of LLMs, researchers have explored integrating LLMs with video encoders to leverage their strong generative and understanding capabilities for video tasks [27, 28, 29]. These studies often utilize open-source LLMs, such as Vicuna [30] and LLaMA [31]. The primary distinction among these works lies in how video features are encoded into visual tokens compatible with LLMs. For example, VideoChat [32] employs vision transformers to encode video features and uses a Query Transformer (Q-Former) [33] to compress video tokens. Similarly, VideoLLaMA [29] combines vision transformers (ViT) and image Q-Formers to encode individual frames, followed by temporal modeling using a video Q-Former. While existing multimodal models perform well on short video tasks, such as captioning or question-answering, they lack fine-grained temporal modeling, rendering them ineffective at capturing deepfake traces in long videos.
@@
Video question-answering datasets. Video question-answering (VQA) is a popular multimodal task, and several large-scale datasets for video-language pretraining have been introduced in recent years. For instance, Ego4D [34] and HowTo100M [35] provide massive internet-scale video data for learning video-language understanding. However, the evaluation objectives of these datasets are often narrow, focusing on specific tasks and lacking a comprehensive assessment of the model’s fine-grained detection capabilities.
@@ III-B Spatio-Temporal Subtle Information Tokenization
As illustrated in Algorithm 1, the ST-SIT comprises two feature extraction branches, feature fusion, and token projection. The first branch introduces the DSEncoder based on the Swin Transformer [36] for local spatio-temporal feature extraction (Lines 1-2). To unify the processing of video frames and explicitly encode temporal information, we propose a 3 × 3 grid layout. Specifically, given a video clip $\mathcal{V}_{\text{clip}}^{(i)}=\{f_{\text{t}}^{(i)}\}_{t=1}^{d}$, we arrange its resized frames $f_{\text{t}}^{\prime(i)}$ into the 3 × 3 as follows:
@@
This grid layout explicitly encodes video temporal dependencies by embedding consecutive frames into spatial neighborhoods. The collection of grid layouts for all video clips is represented as $\mathcal{V}_{\text{grid}}=\{\mathcal{V}_{\text{grid}}^{(1)},\mathcal{V}_{\text{grid}}^{(2)},\dots,\mathcal{V}_{\text{grid}}^{(N)}\}$. Using this unified layout as input for the Swin structure[36], the DSEncoder processes consecutive frames as input and employs a sliding window, combined with hierarchical self-attention mechanisms, to extract spatial details and temporal features. DSEncoder explicitly encodes inter-frame information into spatial neighborhoods through the 3 × 3 compact layout. Additionally, it utilizes a multi-stage window extension mechanism, allowing early layers to focus on capturing local deepfake traces and later layers to emphasize global temporal consistency. The entire encoding process progressively downsamples feature maps through hierarchical patch merging, constructing a multi-scale feature hierarchy. This achieves a dynamic balance between computational efficiency and expressive power, generating compact and robust visual embeddings.
The second branch utilizes the SigLip encoder to extract frame-by-frame visual features (Lines 3-6), leveraging its powerful semantic representation capabilities derived from large-scale vision-language pretraining [37]. To compress SigLip’s frame-by-frame temporal features, we introduce a Compact Visual Connector (CVC). The CVC first utilizes 3D convolution to perform spatio-temporal compression on SigLip’s output features, reducing the number of spatio-temporal tokens to accommodate long video inputs. Then, multiple RegStage [38] blocks are used to supplement and enhance spatio-temporal information, thereby avoiding the loss of critical details. With hierarchical stacking and residual connections, CVC ensures gradient flow and training stability, further enhancing the model’s expressive power and robustness.
The outputs of these two branches, which contain high-dimensional features obtained through deep spatio-temporal modeling and semantic extraction, are integrated via cross-attention [39] to capture their correlation and complementarity (Lines 7-9). This module effectively combines fine-grained spatio-temporal modeling with semantic understanding, incorporating local and global information to construct fused feature representations. Finally, the fused features are projected to tokens via [40] (Lines 10-11) as input to the downstream large language model inference of Fg-MCoT (see next subsection), enabling precise and robust deepfake detection.
@@ III-C Fine-grained Multimodal Chain-of-Thought
which includes the coordinates of the face region, left and right eyes, nose, and the corners of the mouth, along with a confidence score. Based on the coordinate information, we extract only the largest bounding box in each frame and, following the approach of LipForensics[41], add 30% of the face size around the bounding box to refine the bounding box coordinates and write them into $\mathcal{M}_{\text{c}}$. $\mathcal{M}_{\text{c}}$ serves as the supporting preliminary evidence for the next step of facial integrity analysis.
@@
We organize the above metrics into facial integrity metrics $\mathcal{M}_{\Delta}$, including blur variation trends, color distribution consistency, texture feature smoothness, and blending artifacts intensity. $\mathcal{M}_{\Delta}$ is abstracted from raw pixel data, providing clear evidence nodes for downstream rationale generation to suppress hallucinations [17], further constraining the reasoning process to the DVD task. Furthermore, we standardize the data formats of both $\mathcal{M}_{\text{c}}$ and $\mathcal{M}_{\Delta}$ as JSON. This standardized JSON format is designed to facilitate processing by LLM. $\mathcal{M}_{\text{c}}$ and $\mathcal{M}_{\Delta}$ will serve as fine-grained facial information $\mathcal{T}_{\text{fac}}$ input to LLM.
@@
where $\mathcal{T}_{\text{fac}}$ carries highly concise fine-grained semantic representations, ensuring more precise and reliable tracking of forged features. Moreover, facial keypoint pixel localization based on $\mathcal{T}_{\text{fac}}$ can effectively mitigate the inherent hallucination issues in LLM [17].

### v2

@@ Abstract
The rapid development of deepfake video technology has not only facilitated artistic creation but also made it easier to spread misinformation, which is increasingly difficult to identify. Traditional deepfake video detection (DVD) methods face issues such as a lack of transparency in their principles and insufficient generalization capabilities to cope with evolving forgery techniques. This highlights an urgent need for detectors that can identify forged content and provide verifiable reasoning explanations. This paper proposes the explainable deepfake video detection (EDVD) task and designs the EDVD-LLaMA multimodal, a large language model (MLLM) reasoning framework, which provides traceable reasoning processes alongside accurate detection results and trustworthy explanations. Our approach first incorporates a Spatio-Temporal Subtle Information Tokenization (ST-SIT) to extract and fuse global and local cross-frame deepfake features, providing rich spatio-temporal semantic information input for MLLM reasoning. Second, we construct a Fine-grained Multimodal Chain-of-Thought (Fg-MCoT) mechanism, which introduces facial feature data as hard constraints during the reasoning process to achieve pixel-level spatio-temporal video localization, suppress hallucinated outputs, and enhance the reliability of the chain of thought. In addition, we build an Explainable Reasoning FF++ benchmark dataset (ER-FF++set), leveraging structured data to annotate videos and ensure quality control, thereby supporting dual supervision for reasoning and detection. Extensive experiments demonstrate that EDVD-LLaMA achieves outstanding performance and robustness in terms of detection accuracy, explainability, and its ability to handle cross-forgery methods and cross-dataset scenarios. Compared to previous DVD methods, it provides a more explainable and superior solution. The project page is available at: https://11ouo1.github.io/edvd-llama/.
@@ I Introduction
With the increasing popularity of Artificial Intelligence Generated Content (AIGC), deepfake media are easily created using publicly available AIGC video editing tools [av2024latent, skorokhodov2022stylegan, yu2023video]. These techniques can replace faces in videos or alter facial expressions and movements, making such manipulations difficult to detect without a specialized tool [10716705]. While deepfake videos can be used for artistic creation and entertainment, they have serious incidents such as financial fraud, identity theft, and the fabrication of evidence to manipulate public opinion, posing significant information security risks [hondru2025exddv, shao2022open, li2020celeb]. Therefore, developing advanced methods for Deepfake Video Detection (DVD) has become critically important. In this context, the task of DVD aims to determine whether a video is a deepfake. DVD has broad applications in daily work and life, such as verifying the authenticity of individuals in video calls, filtering out fake videos on social media, and preventing the spread of malicious video content.
@@
Recent DVD methods have achieved impressive performance by leveraging carefully designed network architectures, sophisticated constraints, and domain-biased pretraining strategies [10684474, 10620229]. However, existing DVD approaches face two key challenges that limit their generalization and practicality [yan2025generalizing] (see Fig. 2 (a)). First, most DVD methods are black-box models that only output a binary classification result and a confidence score for a video, without revealing the basis of their decisions to users. Because these methods lack interpretable decision rationales, manual re-examination is still required. Second, in the real world, video deepfake techniques are diverse and continuously evolving; existing DVD methods lack comprehensive generalization ability and cannot ensure satisfactory accuracy when confronted with newly emerging deepfake techniques, thereby significantly undermining their practical value.
@@
Given that multimodal large language models (MLLMs) are pretrained on large-scale and diverse corpora of world knowledge, recent studies on explainable image forgery detection [zhang2024common, xu2024fakeshield, huang2024ffaamultimodallargelanguage, zhou2025aigi, Yang_2025_CVPR] have introduced MLLM architectures, which show great potential for enhancing the explainability of detection results and have achieved strong performance. However, these methods share two inherent limitations: first, they operate at the image level and are difficult to apply directly to video, resulting in the loss of temporal features and hindering the recognition of holistic video-level deepfake behaviors. Furthermore, the lack of cross-frame consistency modeling prevents key video deepfake cues, such as boundary artifacts and abnormal physiological rhythms in micro-expressions, from being captured. Second, MLLMs often generate descriptions directly from an image, which frequently produces severe hallucinations that list forged artifacts absent from the sample and undermine the credibility of the generated explanations. In addition, although HEIE [Yang_2025_CVPR] constructs a Chain-of-Thought (CoT) for static images and achieves promising results, its CoT lacks temporal information. It has no structured data as a hard constraint for structured reasoning. This causes its explanations to rely mainly on free-form text, invites semantic hallucination, and limits generalization to complex and temporally sensitive deepfake video recognition.
@@
Motivated by the abovementioned analysis, we propose an Explainable DVD (EDVD) task and a reasoning framework named EDVD-LLaMA, designed to reason and detect deepfake videos while addressing the two significant limitations. As illustrated in Fig. 2 (b), the EDVD task requires models to extract spatio-temporal features from a given video, integrating pixel-level deepfake clues such as edge artifacts and global semantic anomalies like incorrect perspective, to provide clear reasoning principles and ensure reliable accuracy. In the EDVD-LLaMA framework, we first develop a Spatio-Temporal Subtle Information Tokenization (ST-SIT), which comprises the Deepfake Sniffing Encoder (DSEncoder) for extracting local deepfake features. Meanwhile, SigLiP is utilized to extract global video features, followed by the Compact Visual Connector (CVC) for spatio-temporal compression and fusion. Second, we construct a Fine-grained Multimodal Chain-of-Thought (Fg-MCoT) framework based on LLM [qwen2025qwen25technicalreport], which incorporates facial refinement metrics (e.g., landmarks variations) as structured constraints to reduce hallucinations[10.1162/COLI.a.16] and improve reliability. Additionally, we inject structured facial data into the existing deepfake video datasets [roessler2019faceforensicspp] and leverage pretrained models [cheng2024videollama2advancingspatialtemporal] to construct question-video-rationale-answer reasoning quadruples, thereby creating an Explainable Reasoning FF++ Dataset (ER-FF++set). Using this dataset, we further fine-tune our EDVD-LLaMA to equip it with comprehensive analytical capabilities, fine-grained reasoning abilities, and accurate deepfake detection capabilities. Our main contributions are summarized as follows:
@@ II Related Work
Video forgery detection. Yan et al. [Yan_2025_CVPR] propose video-level blending data (VB) and spatiotemporal adapters (StA) to improve the generalization of deepfake video detection, and for the first time simulate the “Facial Feature Drift” forgery artifact. Their method balances spatial and temporal information, but still has room for improvement in handling multimodal forgeries and complex real-world scenarios. The PSO-EfficientNet-GRU method [cunha2024videodeepfake] leverages particle swarm optimization in conjunction with deep learning models to significantly enhance the accuracy of video forgery detection. UFCC [electronics13040804] employs content consistency analysis, offering a unified forensic framework that facilitates more stable identification of tampered regions in images and videos. MSIDSnet[CHENG2024104263] employs multi-scale fusion and a visual Transformer model to achieve robust detection across videos of varying quality. Overall, these methods enhance the performance of video forgery detection from perspectives such as spatiotemporal features, content consistency, multimodal fusion, and model optimization. However, they still face challenges in generalization ability and adaptation to complex real-world environments.
@@
Explainable image forgery detection. Traditional interpretable deepfake detection methods primarily construct explainable networks through model design [trinh2021interpretable] or generation operations [shao2022detecting]. For example, DFGNN [khalid2023dfgnn] applies interpretable GNNs to the deepfake detection task. Although these methods enhance the interpretability of the models, they are still unable to describe the reasoning in natural language. In contrast, recent researchers have utilized MLLMs to construct question-answering tasks [zhang2024common, huang2024ffaamultimodallargelanguage, xu2024fakeshield], providing human-understandable explanations for traditional binary classifiers. For example, CSRDD [zhang2024common] constructs a transferable model based on BLIP, enhancing the generalization ability and interpretability of CNN detectors through multimodal feature fusion. FFAA [huang2024ffaamultimodallargelanguage] redefines deepfake image detection as an open-world visual question answering (VQA) task. They propose a Multi-Answer Intelligent Decision System (MIDS), construct an “image description–forensics–decision” triplet dataset, and utilize LLMs combined with multi-hypothesis prompts to generate reasoning explanations for decisions, demonstrating robustness to unknown forgery techniques. FakeShield [xu2024fakeshield] performs image forgery detection and localization (IFDL) guided by domain labels. A LoRA-tuned MLLM combines tampering classification labels, textual localization rationales, and a segmentation model to achieve synchronous outputs of detection, localization, and supporting evidence, providing interpretable semantic or physical clues and thus improving cross-domain generalization. Although these pioneering works achieve interpretable forgery analysis for single images by integrating visual and language reasoning, they are still limited to static images or facial regions.
@@
Multimodal large language model. With the rapid advancement of LLMs, researchers have explored integrating LLMs with video encoders to leverage their strong generative and understanding capabilities for video tasks [jin2024chatuniviunifiedvisualrepresentation, liu2023videotellerenhancingcrossmodalgeneration, zhang2023videollamainstructiontunedaudiovisuallanguage]. These studies often utilize open-source LLMs, such as Vicuna [NEURIPS2023_91f18a12] and LLaMA [touvron2023llama2openfoundation]. The primary distinction among these works lies in how video features are encoded into visual tokens compatible with LLMs. For example, VideoChat [li2024videochatchatcentricvideounderstanding] employs vision transformers to encode video features and uses a Query Transformer (Q-Former) [li2023blip2bootstrappinglanguageimagepretraining] to compress video tokens. Similarly, VideoLLaMA [zhang2023videollamainstructiontunedaudiovisuallanguage] combines vision transformers (ViT) and image Q-Formers to encode individual frames, followed by temporal modeling using a video Q-Former. While existing multimodal models perform well on short video tasks, such as captioning or question-answering, they lack fine-grained temporal modeling, rendering them ineffective at capturing deepfake traces in long videos.
@@
Video question-answering datasets. Video question-answering (VQA) is a popular multimodal task, and several large-scale datasets for video-language pretraining have been introduced in recent years. For instance, Ego4D [Grauman_2022_CVPR] and HowTo100M [miech2019howto100mlearningtextvideoembedding] provide massive internet-scale video data for learning video-language understanding. However, the evaluation objectives of these datasets are often narrow, focusing on specific tasks and lacking a comprehensive assessment of the model’s fine-grained detection capabilities.
@@ III-B Spatio-Temporal Subtle Information Tokenization
As illustrated in Algorithm 1, the ST-SIT comprises two feature extraction branches, feature fusion, and token projection. The first branch introduces the DSEncoder based on the Swin Transformer [liu2021swintransformerhierarchicalvision] for local spatio-temporal feature extraction (Lines 1-2). To unify the processing of video frames and explicitly encode temporal information, we propose a 3 × 3 grid layout. Specifically, given a video clip $\mathcal{V}_{\text{clip}}^{(i)}=\{f_{\text{t}}^{(i)}\}_{t=1}^{d}$, we arrange its resized frames $f_{\text{t}}^{\prime(i)}$ into the 3 × 3 as follows:
@@
This grid layout explicitly encodes video temporal dependencies by embedding consecutive frames into spatial neighborhoods. The collection of grid layouts for all video clips is represented as $\mathcal{V}_{\text{grid}}=\{\mathcal{V}_{\text{grid}}^{(1)},\mathcal{V}_{\text{grid}}^{(2)},\dots,\mathcal{V}_{\text{grid}}^{(N)}\}$. Using this unified layout as input for the Swin structure[liu2021swintransformerhierarchicalvision], the DSEncoder processes consecutive frames as input and employs a sliding window, combined with hierarchical self-attention mechanisms, to extract spatial details and temporal features. DSEncoder explicitly encodes inter-frame information into spatial neighborhoods through the 3 × 3 compact layout. Additionally, it utilizes a multi-stage window extension mechanism, allowing early layers to focus on capturing local deepfake traces and later layers to emphasize global temporal consistency. The entire encoding process progressively downsamples feature maps through hierarchical patch merging, constructing a multi-scale feature hierarchy. This achieves a dynamic balance between computational efficiency and expressive power, generating compact and robust visual embeddings.
The second branch utilizes the SigLip encoder to extract frame-by-frame visual features (Lines 3-6), leveraging its powerful semantic representation capabilities derived from large-scale vision-language pretraining [zhai2023sigmoidlosslanguageimage]. To compress SigLip’s frame-by-frame temporal features, we introduce a Compact Visual Connector (CVC). The CVC first utilizes 3D convolution to perform spatio-temporal compression on SigLip’s output features, reducing the number of spatio-temporal tokens to accommodate long video inputs. Then, multiple RegStage [radosavovic2020designingnetworkdesignspaces] blocks are used to supplement and enhance spatio-temporal information, thereby avoiding the loss of critical details. With hierarchical stacking and residual connections, CVC ensures gradient flow and training stability, further enhancing the model’s expressive power and robustness.
The outputs of these two branches, which contain high-dimensional features obtained through deep spatio-temporal modeling and semantic extraction, are integrated via cross-attention [vaswani2017attention] to capture their correlation and complementarity (Lines 7-9). This module effectively combines fine-grained spatio-temporal modeling with semantic understanding, incorporating local and global information to construct fused feature representations. Finally, the fused features are projected to tokens via [liu2023visual] (Lines 10-11) as input to the downstream large language model inference of Fg-MCoT (see next subsection), enabling precise and robust deepfake detection.
@@ III-C Fine-grained Multimodal Chain-of-Thought
which includes the coordinates of the face region, left and right eyes, nose, and the corners of the mouth, along with a confidence score. Based on the coordinate information, we extract only the largest bounding box in each frame and, following the approach of LipForensics[haliassos2021lipsdontliegeneralisable], add 30% of the face size around the bounding box to refine the bounding box coordinates and write them into $\mathcal{M}_{\text{c}}$. $\mathcal{M}_{\text{c}}$ serves as the supporting preliminary evidence for the next step of facial integrity analysis.
@@
We organize the above metrics into facial integrity metrics $\mathcal{M}_{\Delta}$, including blur variation trends, color distribution consistency, texture feature smoothness, and blending artifacts intensity. $\mathcal{M}_{\Delta}$ is abstracted from raw pixel data, providing clear evidence nodes for downstream rationale generation to suppress hallucinations [10.1162/COLI.a.16], further constraining the reasoning process to the DVD task. Furthermore, we standardize the data formats of both $\mathcal{M}_{\text{c}}$ and $\mathcal{M}_{\Delta}$ as JSON. This standardized JSON format is designed to facilitate processing by LLM. $\mathcal{M}_{\text{c}}$ and $\mathcal{M}_{\Delta}$ will serve as fine-grained facial information $\mathcal{T}_{\text{fac}}$ input to LLM.
@@
where $\mathcal{T}_{\text{fac}}$ carries highly concise fine-grained semantic representations, ensuring more precise and reliable tracking of forged features. Moreover, facial keypoint pixel localization based on $\mathcal{T}_{\text{fac}}$ can effectively mitigate the inherent hallucination issues in LLM [10.1162/COLI.a.16].